OpenStudy (anonymous):

The height of an equilateral triangular prism increases by one unit. The new lateral area is more than the original by how much? the area of the base the height the perimeter of the base the area of one lateral face

OpenStudy (lgbasallote):

lateral area of triangular prism is 3 x base x height if height will be increased by 1 then you have 3b(h+1) agree?

OpenStudy (anonymous):

so it would be the height

OpenStudy (lgbasallote):

distribute 3b into h+1 what do you get?

OpenStudy (anonymous):

honestly idk

OpenStudy (lgbasallote):

distributive property... \[3b(h+1) = (3b \times h) + (3b \times 1)\] what do you get then?/

OpenStudy (anonymous):

3bh there arent any number to plug in

OpenStudy (amistre64):

what does 3 base signify?

OpenStudy (lgbasallote):

3 x base edge...it was the formule i told her a while ago because it was how she phrased it...i just wanted to be consistent lol

OpenStudy (amistre64):

lateral area is the sum of the side area; each side has an area of: \(s_n*h\) ah, base edge .... got it

OpenStudy (lgbasallote):

@cutie_pie_1213 just 3bh? what about + (3b x 1)?

OpenStudy (anonymous):

ohh okay i thought that part was timesing it by one which u get the same thing

OpenStudy (amistre64):

if we phrase it similar to the choices we get: perimeter * height therefore perimeter(height + 1)

OpenStudy (lgbasallote):

so what's 3b x 1?

OpenStudy (anonymous):

3b

OpenStudy (lgbasallote):

so 3b(h+1) = 3bh = 3b  3bh is LSA (i said this before)...what do you think is 3b?